OpenStudy (anonymous):

Find the value of the square root of 6 over 5 times the square root of 3. in simplest form. 2 over 5. 3 over 4. the square root of 15 over 5. the square root of 2 over 5.

OpenStudy (anonymous):

\[\frac{ \sqrt{6} }{ \sqrt{5} } \times \sqrt{3} = \frac{ \sqrt{6} \times \sqrt{3} }{ \sqrt{5} }\] so this becomes \[\frac{ \sqrt{6 \times 3 } }{\sqrt{5}}\] So solve that and simplify it .